Write the equation of a line that is perpendicular to the given line and that passes through the given point.
–x + 5y = 14; (–5, –2)

Answer by amarjeeth123(569)

The equation of the line is –x + 5y = 14
5y=x+14
y=1/5x+14/5
Slope of the line=1/5
Slope of a line perpendicular to the line=-5
The equation of the line is y=mx+c
y=-5x+c
Since the line passes through (-5,-2)
-2=25+c
c=-27
The equation of the line is y=-5x-27